find the domain of the function
g(x)=2/7-3x

Answer by stanbon(75887)

find the domain of the function
g(x)=2/7-3x
---
7-3x cannot zero.
-------------------
If 7-3x=0
3x = 7
x = 7/3
---------------
So domain is All Real Numbers except x = 7/3
================================================
Cheers,
Stan H.